# Takahiro Nishinakaa,b,c and Yutaka Yoshida d,e

### v1

@@ 4.5 Example: $(A_{2},A_{1})$
where $K_{ab}$ is an $N\times N$ matrix and $B$ is an $N$-dimensional vector. For $N=5$, by choosing $K$ and $B$ as specified below:
@@
we have found experimentally that the resulting Nahm sum agrees with the vacuum character [37] of the $(3,8)$ $W_{3}$ minimal model up to a high order in $q$:
@@
Once the Nahm sum (4.52) is given, it is straightforward to read off the gauge group, CS levels, and chiral multiplet gauge charges of the corresponding CS matter theory. The gauge group is $U(1)^{5}$, the CS level matrix is given by (4.51), and the five chiral multiplets $\{\phi_{i}\}_{i=1}^{5}$ carry diagonal gauge charges $Q_{ai}=\delta_{ai}$ for $a,i=1,\cdots,5$ with the R-charge zero. Without superpotential, the half index with the $(\mathcal{D},D_{c})$ boundary condition is written as
@@
The following linear combination of the topological $U(1)$ symmetries,
$T:=\sum_{a=1}^{4}a\,J^{(a)}+2J^{(5)}\,$, (4.56)
@@
If we take $\eta=q^{-2}$, the half index, which means the R-charge is shifted by $R\to R_{\rm new}=R-2T$ in the definition of the half index, (4.53) reproduces the Nahm sum formula (4.52) for the vacuum character.
@@
$\mathcal{I}^{\rm 4d}(q,t)$ $:=\mathrm{Tr}(-1)^{F}q^{E-R_{\rm 4d}}t^{R_{\rm 4d}-r_{\rm 4d}}\,.$ (4.62)
Here $R_{\rm 4d}$ and $r_{\rm 4d}$ are the generators of 4d $SU(2)\times U(1)$ R-symmetry. Note that, when $q=1$, (4.60) and (4.62) reduce to the H-twisted index and the Schur index, respectively.
@@ 5.1 Schur index for $(A_{3},A_{N-1})$
$\mathcal{I}^{4D}$ $=(q)_{\infty}^{3(N-1)}\sum_{i=1}^{N-1}\sum_{l=1}^{3}\sum_{n^{(l)}_{i}=0}^{\infty}\sum_{k^{(l)}_{i}\in D}\sum_{\widetilde{k}^{(l)}_{i}\in\widetilde{D}}q^{\frac{A+\widetilde{A}}{2}}(-q^{\frac{1}{2}})^{\sum_{i=1}^{N-1}(2\sum_{l=1}^{3}n^{(l)}_{i}-k^{(1)}_{i}-2k^{(2)}_{i}-k^{(3)}_{i}-\widetilde{k}^{(1)}_{i}-2\widetilde{k}^{(2)}_{i}-\widetilde{k}^{(3)}_{i})}$
@@
$\qquad+\sum_{i=1}^{N-1}\Bigl(-\sum_{l=1}^{2}n^{(l)}_{i}n^{(l+1)}_{i}+\sum_{1\leq l<l^{\prime}\leq 3}k^{(l)}_{i}k^{(l^{\prime})}_{i}+(k^{(1)}_{i})^{2}+2(k^{(2)}_{i})^{2}+(k^{(3)}_{i})^{2}\Bigr)\,$, (5.8)
$\widetilde{A}$ $:=A|_{k^{(l)}_{i}\to\widetilde{k}^{(l)}_{i}}$ (5.9)
@@
$D$ $:=\{(k^{(1)}_{i},k^{(2)}_{i},k^{(3)}_{i})\in\mathbb{Z}^{3N}_{\geq 0}|k^{(1)}_{i}+k^{(2)}_{i}\leq n^{(1)}_{i},k^{(1)}_{i}+k^{(2)}_{i}+k^{(3)}_{i}\leq n^{(2)}_{i},k^{(2)}_{i}+k^{(3)}_{i}\leq n^{(3)}_{i}\}$, (5.10)
$\widetilde{D}$ $:=D|_{k^{(l)}_{i}\to\widetilde{k}^{(l)}_{i}}\,$. (5.11)
@@ 5.2 Half index and CS matter theory
and label the 12(N1)12(N-1) chiral multiplets by $\phi^{(l)}_{i},\widetilde{\phi}^{(l)}_{i},\varphi^{(l)}_{i},\widetilde{\varphi}^{(l)}_{i}$ (for $l=1,2,3$ and $i=1,\ldots,N-1$). We assume that the UV (or classical) R-charges of all these chiral multiplets are zero.
@@ 5.3 Monopole superpotential
We consider monopole operators $V^{(l)}_{i}$ and $\widetilde{V}^{(l)}_{i}$ for $l=1,2,3$ and $i=1,\cdots,N-1$, depicted by Table 4. The gauge charge and R-charge is again computed by the formulas (C.1) and (C.2). Then we find that the following dressed monopole operators are gauge invariant:
@@
We define a trial $U(1)$ R-symmetry as the following linear combination of the UV R-charge and the generators of the topological $U(1)$ symmetries:
@@ 5.5 Conjecture for $(A_{M-1},A_{N-1})$
$f(B_{1},T_{1},B_{2},T_{2})$ $\coloneq o(B_{1},T_{1},B_{2},T_{2})-h(B_{1},T_{1},B_{2},T_{2})~,$ (5.69)
$o(B_{1},T_{1},B_{2},T_{2})$ $\coloneq\left\{\begin{array}[]{l}\text{min}(T_{1},T_{2})-\text{max}(B_{1},B_{2})+1\qquad\text{if}\qquad\text{max}(B_{1},B_{2})\leq\text{min}(T_{1},T_{2})\\[5.69054pt] 0\qquad\text{otherwise}\end{array}\right$.~, (5.72)
$h(B_{1},T_{1},B_{2},T_{2})$ $\coloneq\left\{\begin{array}[]{l}\frac{1}{2}\Big(1+(-1)^{o(B_{1},T_{1},B_{2},T_{2})}\Big)\quad\text{if}\quad B_{1}<B_{2}\leq T_{1}<T_{2}\quad\text{or}\quad B_{2}<B_{1}\leq T_{2}<T_{1}\\[5.69054pt] 1\qquad\text{if}\qquad B_{1}=B_{2}\quad\text{or}\quad T_{1}=T_{2}\\[5.69054pt] \frac{1}{2}\Big(1-(-1)^{o(B_{1},T_{1},B_{2},T_{2})}\Big)\quad\text{if}\quad B_{1}<B_{2}<T_{2}<T_{1}\quad\text{or}\quad B_{2}<B_{1}<T_{1}<T_{2}\\[8.53581pt] 0\qquad\text{otherwise}\end{array}\right$.~, (5.77)
$g(B_{1},T_{1},B_{2},T_{2})$ $\coloneq\text{floor}\left(\frac{o(B_{1},T_{1},B_{2},T_{2})}{2}\right)~.$ (5.78)
@@ Appendix A Three dimensional localization formulas
In this appendix we summarize the localization formulas for the half index, the superconformal index, and the ellipsoid partition function of 3d $\mathcal{N}=2$ abelian CS matter theories. Let us consider an abelian CS matter theory with gauge group $U(1)^{N}$ coupled to $L$ chiral multiplets with the gauge charges $Q_{ai}$ for $a=1,\ldots,N$ and $i=1,\ldots,L$ and with $r_{i}$ the R-charge for $i=1,\ldots,L$. We denote the effective gauge CS level [^14] Here we use the term effective in the sense of the half index. In flat space, the level shift usually depends on the sign of the fermion mass, whereas for the half index, the sign of level shift depends on the choice of boundary conditions for the vector and chiral multiplets. This shift was first pointed out in [40] as a regularization factor of the divergence of the one-loop determinants. An interpretation in terms of edge modes was given in [33]. by $K_{ab}$ with $a,b=1,\ldots,N$, and the effective gauge-R-symmetry CS level by $K_{aR}$ with $a,b=1,\ldots,N$. For our purposes, we may set the gauge–flavor mixed CS levels to zero without loss of generality. We will also omit background CS levels, as they contribute an overall multiplicative factor.
@@
Here $F$ is the fermion number, $R$ is the $U(1)$ R-charge, $J_{3}$ is the generator of rotations on the hemisphere $D^{2}$. and $F_{i}$ denotes a $U(1)$ global symmetry charge, which is a topological symmetry acting on monopole operators or a flavor symmetry acting on chiral multiplets. The $\mathcal{H}(D^{2})$ is the space of BPS operators on $D^{2}$.
The half index depends on the choice of boundary conditions for both vector and chiral multiplets at the boundary of $S^{1}\times D^{2}$. For a vector multiplet, there are two standard types of boundary conditions: the Neumann boundary condition ($\mathcal{N}$) [40] and the Dirichlet boundary condition ($\mathcal{D}$)[33]. Since the Dirichlet boundary condition for the vector multiplet is directly related to the Schur index, we will focus on this choice in what follows. For a chiral multiplet, there are again Neumann ($N$) and Dirichlet ($D$) boundary conditions, and in addition one can impose an deformed Dirichlet boundary condition ($D_{c}$). Since a chiral multiplet with the deformed Dirichlet boundary condition acquires a non-zero boundary vev, only those flavor symmetries that leave this vev invariant remain unbroken. Among these, it is the deformed Dirichlet boundary condition that is relevant for the Schur index, and this will be the choice we adopt in the following. Then the half index is written as follows:
@@
where $Q_{i}\cdot n=\sum_{a=1}^{N}Q_{a,i}n_{a}$. $F_{i}$ (resp. $w_{i}$) is the $U(1)$ flavor charge (resp. fugacity) for the $i$-the chiral multiplet. $x_{a}$, $r_{i}$ are same as the above case.
@@ Appendix B Derivation of formula for $\mathcal{I}^{\rm 4d}$ and $\mathcal{S}^{\rm 3d}_{\sf b}$
$\sum_{1\leq\ell<\ell\leq L^{\prime}}A^{\prime}_{\ell\ell^{\prime}}m_{\ell}m_{\ell^{\prime}}:=\sum_{\ell=1}^{\ell^{\prime}-1}\sum_{\ell^{\prime}=2}^{L^{\prime}}P_{k\ell}P_{j\ell}\langle\gamma_{k},\gamma_{j}\rangle m_{\ell}m_{\ell^{\prime}}\,.$ (B.3)
@@
$UPV=\begin{pmatrix}D&{\bf 0}_{2r\times(L^{\prime}-2r)}\end{pmatrix}\,$, (B.6)
@@
Here $(V\cdot p)_{\ell^{\prime}}:=\sum_{\ell=1}^{L^{\prime}}V_{\ell^{\prime}\ell}p_{\ell}$. $s_{\ell},\widetilde{s}_{\ell},x_{\ell}$ and $\widetilde{x}_{\ell}$ for $\ell=1,\cdots,L^{\prime}$ are defined by
$s_{\ell^{\prime}}:=\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}m_{\ell}\,,\quad\widetilde{s}_{\ell^{\prime}}:=\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}\widetilde{m}_{\ell}\,$, (B.10)
$x_{\ell^{\prime}}:=\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}p_{\ell}\,,\quad\widetilde{x}_{\ell^{\prime}}:=\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}\widetilde{p}_{\ell}\,$. (B.11)
We define $\sigma_{a}$ and $Q_{a,i}$ for $a=1,\cdots,2L^{\prime}-2r$ and $i=1,\cdots$
$\sigma_{a}:=\begin{cases}x_{a}&\text{for }a=1,\cdots,L^{\prime},\\[2.0pt] x_{a-L^{\prime}}&\text{for }a=L^{\prime}+1,\cdots,L^{\prime}+2r,\\[2.0pt] \widetilde{x}_{a-L^{\prime}-2r}&\text{if }a=L^{\prime}+2r+1,\cdots,2L^{\prime}-2r\,.\end{cases}$ (B.12)
@@
$n_{a}:=\begin{cases}s_{a}&\text{for }a=1,\cdots,L^{\prime},\\[2.0pt] s_{a-L^{\prime}}&\text{for }a=L^{\prime}+1,\cdots,L^{\prime}+2r,\\[2.0pt] \widetilde{s}_{a-L^{\prime}-2r}&\text{if }a=L^{\prime}+2r+1,\cdots,2L^{\prime}-2r\,.\end{cases}$ (B.13)
@@
$Q^{T}_{i,a}:=\begin{cases}V_{i,a}&\text{for }i=1,\cdots L^{\prime},a=1,\cdots L^{\prime},\\[2.0pt] V_{i-L^{\prime},a}&\text{for }i=L^{\prime}+1,\cdots,2L^{\prime},a=1,\cdots,2r\,,\\[2.0pt] V_{a-L^{\prime},i-L^{\prime}+2r}&\text{for }i=L^{\prime}+1,\cdots,2L^{\prime},a=L^{\prime}+1,\cdots 2L^{\prime}-2r\,,\\[2.0pt] 0&\text{the others}\,.\end{cases}$ (B.14)
@@
$\mathcal{I}^{\text{4d}}$ $=(q)^{L-N}_{\infty}\sum_{n_{1},\cdots,n_{N}=0}^{\infty}\,\sum_{(n\cdot Q)_{1}\geq 0,\cdots(n\cdot Q)_{L}\geq 0}\frac{q^{\frac{1}{2}\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}}(-q^{\frac{1}{2}})^{\sum_{a=1}^{N}\sum_{i=1}^{L}Q_{a,i}n_{a}}}{\prod_{i=1}^{L}(q)_{\sum_{a=1}^{N}n_{a}Q_{a,i}}}\,$, (B.15)
@@
$L:=2L^{\prime}\,,\quad N:=2L^{\prime}-2r\,$, (B.17)
and $(n\cdot Q)_{i}:=\sum_{a=1}^{N}n_{a}Q_{ai}$. $K_{ab}$ for $a,b=1,\cdots,N$ is defined by the following relation:
$\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}:=\sum_{\ell=1}^{\ell^{\prime}-1}\sum_{\ell^{\prime}=2}^{L^{\prime}}P_{k\ell}P_{j{\ell}^{\prime}}\langle\gamma_{k},\gamma_{j}\rangle\Bigl((n\cdot Q)_{\ell}(n\cdot Q)_{\ell^{\prime}}+(n\cdot Q)_{\ell+L^{\prime}}(n\cdot Q)_{\ell^{\prime}+L^{\prime}}\Bigr)\,.$ (B.18)

### v2

@@ 4.5 Example: $(A_{2},A_{1})$
where $K_{ab}$ is an $N\times N$ matrix and $B_{a}$ is an $N$-dimensional vector. For $N=5$, by choosing $K$ and $B$ as specified below:
@@
we have found experimentally that the resulting Nahm sum agrees with the vacuum character [37] of the $(3,8)$ $W_{3}$ minimal model up to a high order in qq: [^13] After uploading the first version to arXiv, we are informed by S. O. Warnaar that (4.50) can be derived from Eq. (7.6) of [38]. We are grateful to him for this remark.
@@
Once the Nahm sum (4.50) is given, it is straightforward to read off the gauge group, CS levels, and chiral multiplet gauge charges of the corresponding CS matter theory. The gauge group is $U(1)^{5}$, the CS level matrix is given by (4.49), and the five chiral multiplets $\{\phi_{i}\}_{i=1}^{5}$ carry diagonal gauge charges $Q_{a,i}=\delta_{ai}$ for $a,i=1,\cdots,5$ with the R-charge zero. Without superpotential, the half index with the $(\mathcal{D},D_{c})$ boundary condition is written as
@@
The following linear combination of the topological $U(1)$ symmetries:
$T\equiv\sum_{a=1}^{4}a\,J^{(a)}+2J^{(5)}\,$, (4.54)
@@
If we take $\eta=q^{-2}$, the half index, which means the R-charge is shifted by $R\to R_{\rm shift}=R-2T$ in the definition of the half index, (4.51) reproduces the Nahm sum formula (4.50) for the vacuum character.
@@
$\mathcal{I}^{\rm 4d}(q,t)$ $\equiv\mathrm{Tr}(-1)^{F}q^{E-R_{\rm 4d}}t^{R_{\rm 4d}-r_{\rm 4d}}\,.$ (4.60)
Here $R_{\rm 4d}$ and $r_{\rm 4d}$ are the generators of 4d $SU(2)\times U(1)$ R-symmetry. Note that, when $t=1$, (4.58) and (4.60) reduce to the H-twisted index and the Schur index, respectively.
@@ 5.1 Schur index for $(A_{3},A_{N-1})$
$\mathcal{I}^{\rm 4d}$ $=(q)_{\infty}^{3(N-1)}\sum_{i=1}^{N-1}\sum_{l=1}^{3}\sum_{n^{(l)}_{i}=0}^{\infty}\sum_{k^{(l)}_{i}\in D}\sum_{\widetilde{k}^{(l)}_{i}\in\widetilde{D}}q^{\frac{A+\widetilde{A}}{2}}(-q^{\frac{1}{2}})^{\sum_{i=1}^{N-1}(2\sum_{l=1}^{3}n^{(l)}_{i}-k^{(1)}_{i}-2k^{(2)}_{i}-k^{(3)}_{i}-\widetilde{k}^{(1)}_{i}-2\widetilde{k}^{(2)}_{i}-\widetilde{k}^{(3)}_{i})}$
@@
$\qquad++\sum_{i=1}^{N-1}\Bigl(-\sum_{l=1}^{2}n^{(l)}_{i}n^{(l+1)}_{i}+\sum_{1\leq l<l^{\prime}\leq 3}k^{(l)}_{i}k^{(l^{\prime})}_{i}+(k^{(1)}_{i})^{2}+2(k^{(2)}_{i})^{2}+(k^{(3)}_{i})^{2}\Bigr)\,$, (5.8)
$\widetilde{A}$ $\equiv A|_{k^{(l)}_{i}\to\widetilde{k}^{(l)}_{i}}\,$, (5.9)
@@
$D$ $\equiv\{(k^{(1)}_{i},k^{(2)}_{i},k^{(3)}_{i})\in\mathbb{Z}^{3N}_{\geq 0}|k^{(1)}_{i}+k^{(2)}_{i}\leq n^{(1)}_{i},k^{(1)}_{i}+k^{(2)}_{i}+k^{(3)}_{i}\leq n^{(2)}_{i},k^{(2)}_{i}+k^{(3)}_{i}\leq n^{(3)}_{i}\}$, (5.10)
$\widetilde{D}$ $\equiv D|_{k^{(l)}_{i}\to\widetilde{k}^{(l)}_{i}}\,$. (5.11)
@@ 5.2 Half index and CS matter theory
and label the 12(N1)12(N-1) chiral multiplets by $\phi^{(l)}_{i},\widetilde{\phi}^{(l)}_{i},\varphi^{(l)}_{i},\widetilde{\varphi}^{(l)}_{i}$ (for $l=1,2,3$ and $i=1,\ldots,N-1$). We assume that the UV (or reference) R-charges of all these chiral multiplets are zero.
@@ 5.3 Monopole superpotential
We consider monopole operators $V^{(l)}_{i}$ and $\widetilde{V}^{(l)}_{i}$ for $l=1,2,3$ and $i=1,\cdots,N-1$, depicted by Table 4. The gauge charge and R-charge are again computed by the formulas (C.1) and (C.2). Then we find that the following dressed monopole operators are gauge invariant:
@@
We define a shifted $U(1)$ R-symmetry as the following linear combination of the reference R-charge and the generators of the topological $U(1)$ symmetries:
@@ 5.5 Conjecture for $(A_{M-1},A_{N-1})$
$f(B_{1},T_{1},B_{2},T_{2})$ $\equiv o(B_{1},T_{1},B_{2},T_{2})-h(B_{1},T_{1},B_{2},T_{2})~,$ (5.69)
$o(B_{1},T_{1},B_{2},T_{2})$ $\equiv\left\{\begin{array}[]{l}\text{min}(T_{1},T_{2})-\text{max}(B_{1},B_{2})+1\qquad\text{if}\qquad\text{max}(B_{1},B_{2})\leq\text{min}(T_{1},T_{2})\\[5.69054pt] 0\qquad\text{otherwise}\end{array}\right$.~, (5.72)
$h(B_{1},T_{1},B_{2},T_{2})$ $\equiv\left\{\begin{array}[]{l}\frac{1}{2}\Big(1+(-1)^{o(B_{1},T_{1},B_{2},T_{2})}\Big)\quad\text{if}\quad B_{1}<B_{2}\leq T_{1}<T_{2}\quad\text{or}\quad B_{2}<B_{1}\leq T_{2}<T_{1}\\[5.69054pt] 1\qquad\text{if}\qquad B_{1}=B_{2}\quad\text{or}\quad T_{1}=T_{2}\\[5.69054pt] \frac{1}{2}\Big(1-(-1)^{o(B_{1},T_{1},B_{2},T_{2})}\Big)\quad\text{if}\quad B_{1}<B_{2}<T_{2}<T_{1}\quad\text{or}\quad B_{2}<B_{1}<T_{1}<T_{2}\\[8.53581pt] 0\qquad\text{otherwise}\end{array}\right$.~, (5.77)
$g(B_{1},T_{1},B_{2},T_{2})$ $\equiv\text{floor}\left(\frac{o(B_{1},T_{1},B_{2},T_{2})}{2}\right)~.$ (5.78)
@@ Appendix A Three dimensional localization formulas
In this appendix we summarize the localization formulas for the half index, the superconformal index, and the ellipsoid partition function of 3d $\mathcal{N}=2$ abelian CS matter theories. Let us consider an abelian CS matter theory with gauge group $U(1)^{N}$ coupled to $L$ chiral multiplets with the gauge charges $Q_{a,i}$ for $a=1,\ldots,N$ and $i=1,\ldots,L$ and with the R-charge $r_{i}$ for $i=1,\ldots,L$. We denote the effective gauge CS level [^15] Here we use the term effective in the sense of the half index. In flat space, the level shift usually depends on the sign of the fermion mass, whereas for the half index, the sign of level shift depends on the choice of boundary conditions for the vector and chiral multiplets. This shift was first pointed out in [41] as a regularization factor of the divergence of the one-loop determinants. An interpretation in terms of edge modes was given in [33]. by $K_{ab}$ with $a,b=1,\ldots,N$, and the effective gauge-R-symmetry CS level by $K_{aR}$ with $a,b=1,\ldots,N$. For our purposes, we may set the gauge–flavor mixed CS levels to zero without loss of generality. We will also omit background CS levels, as they contribute an overall multiplicative factor.
@@
Here $F$ is the fermion number, $R$ is the $U(1)$ R-charge, $J_{3}$ is the generator of rotations on the hemisphere $D^{2}$, and $F_{i}$ denotes a $U(1)$ global symmetry charge, which is a topological symmetry acting on monopole operators or a flavor symmetry acting on chiral multiplets. The $\mathcal{H}(D^{2})$ is the space of BPS operators on $D^{2}$.
The half index depends on the choice of boundary conditions for both vector and chiral multiplets at the boundary of $S^{1}\times D^{2}$. For a vector multiplet, there are two standard types of boundary conditions: the Neumann boundary condition ($\mathcal{N}$) [41] and the Dirichlet boundary condition ($\mathcal{D}$)[33]. Since the Dirichlet boundary condition for the vector multiplet is directly related to the Schur index, we will focus on this choice in what follows. For a chiral multiplet, there are again Neumann ($N$) and Dirichlet ($D$) boundary conditions, and in addition one can impose a deformed Dirichlet boundary condition ($D_{c}$). Since a chiral multiplet with the deformed Dirichlet boundary condition acquires a non-zero boundary vev, only those flavor symmetries that leave this vev invariant remain unbroken. Among these, it is the deformed Dirichlet boundary condition that is relevant for the Schur index, and this will be the choice we adopt in the following. Then the half index is written as follows:
@@
where $Q_{i}\cdot n=\sum_{a=1}^{N}Q_{a,i}n_{a}$. $F_{i}$ (resp. $w_{i}$) is the $U(1)$ flavor charge (resp. fugacity) for the $i$-th chiral multiplet. $x_{a}$, $r_{i}$ are same as the above case.
@@ Appendix B Derivation of formula for $\mathcal{I}^{\rm 4d}$ and $\mathcal{S}^{\rm 3d}_{\sf b}$
$\sum_{1\leq\ell<\ell\leq L^{\prime}}A^{\prime}_{\ell\ell^{\prime}}m_{\ell}m_{\ell^{\prime}}\equiv\sum_{\ell=1}^{\ell^{\prime}-1}\sum_{\ell^{\prime}=2}^{L^{\prime}}P_{k\ell}P_{j\ell}\langle\gamma_{k},\gamma_{j}\rangle m_{\ell}m_{\ell^{\prime}}\,.$ (B.3)
@@
$UPV=\begin{pmatrix}D&{\bf 0}_{2r\times(L^{\prime}-2r)}\\ \end{pmatrix}\,$, (B.6)
@@
Here $(V\cdot p)_{\ell^{\prime}}\equiv\sum_{\ell=1}^{L^{\prime}}V_{\ell^{\prime}\ell}p_{\ell}$. $s_{\ell},\widetilde{s}_{\ell},x_{\ell}$ and $\widetilde{x}_{\ell}$ for $\ell=1,\cdots,L^{\prime}$ are defined by
$s_{\ell^{\prime}}\equiv\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}m_{\ell}\,,\quad\widetilde{s}_{\ell^{\prime}}\equiv\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}\widetilde{m}_{\ell}\,$, (B.10)
$x_{\ell^{\prime}}\equiv\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}p_{\ell}\,,\quad\widetilde{x}_{\ell^{\prime}}\equiv\sum_{\ell=1}^{L^{\prime}}(V^{-1})_{\ell^{\prime}\ell}\widetilde{p}_{\ell}\,$. (B.11)
We define $\sigma_{a}$ and $Q_{a,i}$ for $a=1,\cdots,2L^{\prime}-2r$ and $i=1,\cdots,2L^{\prime}$ by
$\sigma_{a}\equiv\begin{cases}x_{a}&\text{for }a=1,\cdots,L^{\prime},\\[2.0pt] x_{a-L^{\prime}}&\text{for }a=L^{\prime}+1,\cdots,L^{\prime}+2r,\\[2.0pt] \widetilde{x}_{a-L^{\prime}-2r}&\text{if }a=L^{\prime}+2r+1,\cdots,2L^{\prime}-2r\,.\end{cases}$ (B.12)
@@
$n_{a}\equiv\begin{cases}s_{a}&\text{for }a=1,\cdots,L^{\prime},\\[2.0pt] s_{a-L^{\prime}}&\text{for }a=L^{\prime}+1,\cdots,L^{\prime}+2r,\\[2.0pt] \widetilde{s}_{a-L^{\prime}-2r}&\text{if }a=L^{\prime}+2r+1,\cdots,2L^{\prime}-2r\,.\end{cases}$ (B.13)
@@
$Q_{a,i}\equiv\begin{cases}V_{i,a}&\text{for }i=1,\cdots L^{\prime},a=1,\cdots L^{\prime},\\[2.0pt] V_{i-L^{\prime},a}&\text{for }i=L^{\prime}+1,\cdots,2L^{\prime},a=1,\cdots,2r\,,\\[2.0pt] V_{a-L^{\prime},i-L^{\prime}+2r}&\text{for }i=L^{\prime}+1,\cdots,2L^{\prime},a=L^{\prime}+1,\cdots 2L^{\prime}-2r\,,\\[2.0pt] 0&\text{the others}\,.\end{cases}$ (B.14)
@@
$\mathcal{I}^{\text{4d}}$ $=(q)^{L-N}_{\infty}\sum_{n_{1},\cdots,n_{N}\in\mathbb{Z}\atop(n\cdot Q)_{1}\geq 0,\cdots(n\cdot Q)_{L}\geq 0}\frac{q^{\frac{1}{2}\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}}(-q^{\frac{1}{2}})^{\sum_{a=1}^{N}\sum_{i=1}^{L}Q_{a,i}n_{a}}}{\prod_{i=1}^{L}(q)_{\sum_{a=1}^{N}n_{a}Q_{a,i}}}\,$, (B.15)
@@
$L\equiv 2L^{\prime}\,,\quad N\equiv 2L^{\prime}-2r\,$, (B.17)
and $(n\cdot Q)_{i}\equiv\sum_{a=1}^{N}n_{a}Q_{ai}$. $K_{ab}$ for $a,b=1,\cdots,N$ is defined by the following relation:
$\sum_{a,b=1}^{N}K_{ab}n_{a}n_{b}\equiv\sum_{\ell=1}^{\ell^{\prime}-1}\sum_{\ell^{\prime}=2}^{L^{\prime}}P_{k\ell}P_{j{\ell}^{\prime}}\langle\gamma_{k},\gamma_{j}\rangle\Bigl((n\cdot Q)_{\ell}(n\cdot Q)_{\ell^{\prime}}+(n\cdot Q)_{\ell+L^{\prime}}(n\cdot Q)_{\ell^{\prime}+L^{\prime}}\Bigr)\,.$ (B.18)